OpenStudy (anonymous):

Show sin^2 x cos^2 x = (1 – cos(4x)) / 8 by filling in the table below. Start with the left hand side.

OpenStudy (anonymous):



OpenStudy (anonymous):

@RadEn this is the last proof for today, i promise :)

OpenStudy (raden):

sin^2 x = (1-cos(2x))/2 and cos^2 x = (1+cos(2x))/2 reason = those are identities the simplify form would be 1/4 (1 - cos^2 (2x)) cos^2 (2x) = (1 + cos(4x))/2 reason = an identity like the first reason 1/4 (1 - cos^2 (2x)) = 1/4 ( 1 - (1 + cos(4x))/2 ) = 1/8(1 - cos(4x)) reason = just simplifying form the 2nd reason maybe :)

OpenStudy (anonymous):

haha thanks, but what is the maybe about?

OpenStudy (raden):

maybe my reason is right, or it can be wrong too :) this link might useful : http://2000clicks.com/mathhelp/GeometryTrigEquivIdentities00.aspx

OpenStudy (anonymous):

Ok ok, but do you think it is right? Like it does make sense haha so I don't see why it would be completely wrong

OpenStudy (raden):

i hope my reason is right :)

OpenStudy (anonymous):

what identities are used in the reasons?

OpenStudy (anonymous):

@RadEn